Solve the equation for the indicated variable. (Enter your answers as a comma-separated list.) ax + b 2; for x cx + d X-
Answer #1

given, (ax+b)/(cx+d)=2

or,  (ax+b)=2(cx+d)

or,  ax+b=2cx+2d

or , ax-2cx=2d-b

or,  x(a-2c)=(2d-b)

or, x=(2d-b)/(a-2c)